OpenStudy (anonymous):

change each of teh following from exponential form to logarithmic form. a)6^-2=1/36 b)m^w=p

jimthompson5910 (jim_thompson5910):

a) \[\Large 6^{-2}=\frac{1}{36}\] \[\Large \log_{6}\left(\frac{1}{36}\right)=-2\] ------------------------------------------------------- b) \[\Large m^w=p\] \[\Large \log_{m}\left(p\right)=w\]